Logiweb dictionary of factorial

Up Help

0 0 factorial
1 2 nani setu kunu...:1

2 1 ∗!

The pyk compiler, version 0.grue.20050502 by Klaus Grue GRD-2005-06-02.UTC:14:20:27.677282 = MJD-53523.TAI:14:20:59.677282 = LGT-4624438859677282e-6